# HAPO: Training Language Models to Reason Concisely via History-Aware Policy Optimization

### v1

@@ 5.2 Comparison with Prior Work
HAPO achieves a superior length-accuracy trade-off compared with prior work. We compare HAPO against L1-Exact, L1-Max, and Query-Opt. On MATH500 and AIME2024, HAPO attains comparable accuracy to the baselines while producing the shortest responses. On GSM8K, HAPO is less effective than Query-Opt and generates longer responses, though it still outperforms L1-Exact and L1-Max. On average, HAPO’s responses are shorter than those of L1-Exact, L1-Max, and Query-Opt by 23%, 20%, and 19%, respectively, while maintaining similar accuracy.
5.3 Case Study on MATH500
@@
5.4 Results on Out of Domain Benchmarks
Table 3: Results on GQPA and LiveCodeBench.
GPQA CodeBench
Models Passe@1 #Tokens Pass@1 #Tokens
DeepSeek-R1-1.5B 34.34 10296 16.42 15070
L1-Exact (U) 36.70 6658 16.23 10965
L1-Max (U) 33.43 6168 16.01 10578
Query-Opt (Q) 35.86 8228 17.05 11456
HAPO w/ $w=0$ 33.33 8611 16.67 11098
HAPO 32.32 7563 16.84 9933
We evaluate HAPO on two out-of-domain benchmarks to assess its generalizability. GPQA [25] comprises multiple-choice questions on biology, physics, and chemistry, while LiveCodeBench [16] includes a variety of coding-related problems.
As shown in Table 3, training on math problems with HAPO enables the model to reason more concisely in other domains, albeit to a lesser extent. On GPQA, HAPO reduces response length by 27% with a 2% drop in accuracy. On LiveCodeBench, it achieves a 34% reduction in length while slightly improving accuracy. Compared to Query-Opt, HAPO produces shorter responses but with lower accuracy on both benchmarks. As for the L1 methods, HAPO outperforms them on LiveCodeBench but underperforms on GPQA. Further investigation into these domain-specific trade-offs is left for future work.
5.5 Training Dynamics
@@
5.6 Ablation Study on Length Reward
We conduct an ablation study on the length reward weight constant $w$, and present the results in Table 4. The comparison between $w=1$ and $w=0$ aligns with expectations: setting $w=0$, where the model is trained solely with the correctness reward, results in higher accuracy but longer responses.
Table 4: Results for the ablation study.
GSM8K MATH500 AIME2024 Average
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO ($w=1$) 79.08 661 81.05 2978 26.56 8171 62.23 3937
w/ $w=0.5$ 74.54 370 79.80 4057 24.44 7827 59.59 4085
w/ $w=0.1$ 74.94 473 77.35 2910 24.79 8418 59.03 3934
w/ $w=0$ 84.40 1381 82.15 4939 31.15 10701 66.02 5674
Somewhat counterintuitively, using intermediate values such as $w=0.1$ or $w=0.5$ leads to lower accuracy compared to $w=1$. While response length fluctuates on MATH500 and AIME2024, it decreases on GSM8K. We suspect that this behavior may be attributed to the internal design of GRPO, which normalizes the final reward using the mean and standard deviation of the rewards within each batch. This normalization could partially offset the intended effect of adjusting $w$. We leave further investigation of this phenomenon to future work.
5.7 Training on Smaller Datasets for More Epochs
Prior works such as [8] suggest that training LLMs on a small set of occasionally solvable examples can effectively reduce their response length. Inspired by this, we check whether reducing the size of our training set will yield similar or even better performance. In particular, we introduce two new settings, where we reduce the training set size from the original 2,000 to 500 and 100 respectively. In each setting, we train the model for 10 epochs to ensure a sufficient number of iterations while preventing overfitting. We keep other hyperparameters the same. Table 5 shows the results.
GSM8K MATH500 AIME2024 Average
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO ($|\mathcal{D}|=2$k) 79.08 661 81.05 2978 26.56 8171 62.23 3937
w/ $|\mathcal{D}|=500$ 84.06 1371 82.05 5194 29.90 10700 65.34 5755
w/ $|\mathcal{D}|=100$ 83.89 1785 80.80 5887 30.94 13518 65.21 7063
Compared with the original setting, training on fewer examples leads to higher accuracy but longer responses. The overall performance becomes much closer to that of the base model. This is expected: with a smaller training set, the model is exposed to a more limited range of question types, increasing the likelihood of encountering unfamiliar questions during evaluation. As discussed in Section 5.4, the concise reasoning abilities learned by HAPO do generalize to unseen questions, but to a limited extent. Consequently, the model produces longer responses and deviates less from the base model.
6 Conclusion
In conclusion, we propose History-Aware Policy Optimization (HAPO), a training method that enables large language models to reason more concisely by leveraging historical information from past interactions. HAPO encourages models to iteratively discover correct responses that are more concise than those previously found. This mechanism leads to solutions that are both accurate and cost-efficient at inference time. Experimental results demonstrate that HAPO significantly reduces response length while maintaining accuracy, achieving a more favorable length-correctness trade-off than prior approaches. Moreover, HAPO generalizes effectively to questions from diverse sources and difficulty levels. Through the use of training history, HAPO opens new directions for building more adaptive, efficient, and capable models in reasoning tasks.
Appendix A Visualization of Results for DeepSeek-R1-1.5B
@@ 6 Conclusion
Appendix B Results with PPO
In addition to GRPO, we also experiment with PPO [26] as the training algorithm. We set the learning rate to 3e-6 and train the models for 5 epochs. The results obtained using PPO are presented in Table 6.
Table 6: PPO results on GSM8K, MATH500 and AIME2024.
@@ Appendix B Results with PPO
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
@@
Appendix C Alternative History State
It is also possible to update $h_{i}$ using the mean function instead of minimum, allowing $h_{i}$ to track the average length of correct responses in the training history. In this case, the model receives a positive length reward whenever it generates a correct response that is shorter than the average of the previous correct ones.
Table 7: Results comparing different types of $h_{i}$.
GSM8K MATH500 AIME2024 Average
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO w/ $h_{i}=min$ 79.08 661 81.05 2978 26.56 8171 62.23 3937
HAPO w/ $h_{i}=mean$ 74.46 413 74.90 2676 18.33 7487 55.90 3523
Under this new update schedule, we expect $h_{i}$ to decrease more slowly during training, implying a slower reduction in response length. However, the results in Table 7 suggest otherwise. With the same number of training epochs, HAPO with the mean update produces more concise responses but lower accuracy than HAPO with the min update across all benchmarks. This indicates that HAPO tends to reduce response length more aggressively when using the historical mean. A possible explanation lies in the reward normalization step of GRPO. When the min function is used, the lengths of most of the newly sampled responses are likely to exceed $h_{i}$, making the length rewards concentrate in the negative range—resulting in lower reward variance and weaker learning signals. In contrast, using the mean function allows the sampled lengths to be more symmetrically distributed around $h_{i}$, leading to higher reward variance and thus stronger learning signals.
Appendix D Repetition Reward
Following Yeo et al. [45], we incorporate an additional repetition penalty into the reward function. Specifically, we identify repeated n-grams within each generated response, count the total number of tokens that appear in these repeated n-grams, and divide this count by the total number of tokens in the response. The resulting repetition penalty, denoted as $rp$, is scaled by a weight constant $w_{\text{rp}}$ and subtracted from the overall reward. The final reward is given by:
Table 8: Results with repetition reward.
GSM8K MATH500 AIME2024 Average
Models Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens Pass@1 #Tokens
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO 79.08 661 81.05 2978 26.56 8171 62.23 3937
w/ Repetition Penalty 73.57 404 73.05 2782 17.71 6996 54.78 3394
Compared to HAPO, incorporating the repetition penalty results in shorter responses but leads to additional accuracy drops of approximately 5%, 8%, and 9% on GSM8K, MATH500, and AIME2024, respectively. We suspect that the penalty may be too strong. Using larger n-grams and reducing the weight $w_{\text{rp}}$ could potentially improve overall performance.
Appendix E Discussion
E.1 Limitations and Future Work
@@ E.2 Broader Impact
Appendix F Training and Reproduction Details
We train our models using the Adam optimizer with a learning rate of 1e-6 and a linear learning rate scheduler. The batch size is set to 2, with a gradient accumulation step of 4. During training, we sample 2 responses from $p_{\theta}$ for each query in a batch using a temperature of 0.7. The best checkpoint is selected based on the performance on a separate set of 500 examples from the DeepScaleR-Preview-Dataset. All experiments are conducted using 4 A6000 GPUs and 2 A100 GPUs.
In both training and evaluation, we use the following system prompt
[…]
and set the user message to the question text to get the model response.

### v2

@@ Introduction
HAPO achieves a superior length-accuracy trade-off compared with prior work. In Table 2, we compare HAPO against L1-Exact, L1-Max, Query-Opt, and O1-Pruner.
On MATH500, HAPO produces the shortest responses, using 9% fewer tokens than the second place. It attains an accuracy that is similar to the prior methods, except for O1-Pruner, whose accuracy is much lower. On AIME2024, HAPO again gives the most concise responses, using 19% fewer tokens than the second place. In terms of accuracy, it is superior or on par with prior methods, except that O1-Pruner’s accuracy stands out higher. On GSM8K, HAPO is less effective than Query-Opt and generates longer responses, though it still outperforms the other methods.
On average, HAPO’s responses are shorter than L1-Exact, L1-Max, Query-Opt, and O1-Pruner by 23%, 20%, 19%, and 16% respectively, while maintaining similar accuracy.
Case Study on MATH500
@@
Results on Out of Domain Benchmarks
GPQA CodeBench Average
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$
DeepSeek-R1-1.5B 34.34 10296 16.42 15070 25.38 12683
L1-Exact (U) 36.70 6658 16.23 10965 26.47 8812
L1-Max (U) 33.43 6168 16.01 10578 24.72 8373
Query-Opt (Q) 35.86 8228 17.05 11456 26.46 9842
O1-Pruner (Q) 34.85 7944 16.34 10609 25.60 9277
HAPO w/ $w=0$ 33.33 8611 16.67 11098 25.00 9855
HAPO 32.32 7563 16.84 9933 24.58 8748
Table 3: Results on GPQA and LiveCodeBench.
We evaluate HAPO on two out-of-domain benchmarks to assess its generalizability. GPQA (rein2024gpqa) comprises multiple-choice questions on biology, physics, and chemistry, while LiveCodeBench (jain2025) includes a variety of coding-related problems.
GSM8K MATH500 AIME2024 Average
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO ($|\mathcal{D}|=2$k) 79.08 661 81.05 2978 26.56 8171 62.23 3937
w/ $|\mathcal{D}|=1k$ 82.70 1108 83.55 3463 25.00 10408 63.75 4993
w/ $|\mathcal{D}|=500$ 84.06 1371 82.05 5194 29.90 10700 65.34 5755
w/ $|\mathcal{D}|=100$ 83.89 1785 80.80 5887 30.94 13518 65.21 7063
As shown in Table 3, training on math problems with HAPO enables the model to reason more concisely in other domains, albeit to a lesser extent. On GPQA, HAPO reduces response length by 27% with a 2% drop in accuracy. On LiveCodeBench, it achieves a 34% reduction in length while slightly improving accuracy.
Compared with prior work, HAPO achieves the best length-correctness tradeoffs on LiveCodeBench, with the shortest responses and similar accuracy. On GPQA, the results are mixed. HAPO produces shorter responses but lower accuracy than Query-Opt and O1-Pruner, and it underperforms the L1 methods. Further investigation into these domain-specific trade-offs is left for future work.
Training Dynamics
@@
Training on Smaller Datasets for More Epochs
Prior works such as (fatemi2025concise) suggest that training LLMs on a small set of occasionally solvable examples can effectively reduce their response length. Inspired by this, we check whether reducing the size of our training set will yield similar or even better performance. In particular, we reduce the training set size from the original 2,000 to 1,000, 500, and 100, respectively. In each setting, we train the model for 10 epochs to ensure a sufficient number of iterations while preventing overfitting. We keep other hyperparameters the same. Table 4 shows the results.
Compared with the original setting, training on fewer examples leads to higher accuracy but longer responses. The overall performance becomes much closer to that of the base model. This is expected: with a smaller training set, the model is exposed to a more limited range of question types, increasing the likelihood of encountering unfamiliar questions during evaluation. As discussed previously, the concise reasoning abilities learned by HAPO do generalize to unseen questions, but to a limited extent. Consequently, the model deviates less from the base model.
Conclusion
In conclusion, we propose History-Aware Policy Optimization (HAPO), a training method that enables large language models to reason more concisely by leveraging historical information from past interactions. HAPO encourages models to iteratively discover correct responses that are more concise than those previously found. Experimental results demonstrate that HAPO significantly reduces response length while maintaining accuracy, achieving a more favorable length-correctness trade-off than prior approaches. Through the use of training history, HAPO opens new directions for building more adaptive, efficient, and capable models in reasoning tasks.
Visualization of Results for DeepSeek-R1-1.5B
@@ Conclusion
Reward Design
Our reward design ensures that correct responses always receive strictly higher overall rewards than incorrect responses. In particular, the overall reward of any correct response is within the range $[1+wc,1+w]$, where the lower bound $1+wc$ is achieved when $|y_{i}|\geq 2h_{i}$. Note than $c\in(-1,0)$ and $w\in[0,1]$, so cw>1cw>-1 and $1+wc>0$. The overall reward of any incorrect response is within the range [w,0][-w,0], where the upper bound 0 is achieved when $|y_{i}|\leq h_{i}$.
Training and Reproduction Details
We train our models using the Adam optimizer with a learning rate of 1e-6 and a linear learning rate scheduler. The batch size is set to 2, with a gradient accumulation step of 4. During training, we sample responses from $p_{\theta}$ for each query in a batch using a temperature of 0.7. The best checkpoint is selected based on the performance on a separate set of 500 examples from the DeepScaleR-Preview-Dataset. All experiments are conducted using 4 A6000 GPUs and 2 A100 GPUs.
In both training and evaluation, we use the following system prompt
[…]
and set the user message to the question text to get the model’s response.
Study on History State
We experiment with different choices of $h_{i}$ to investigate its impact on training. First, to experiment with other possible designs of $h_{i}$, we change the aggregation function from minimum to mean, in which case $h_{i}$ tracks the average length of correct responses in the training history. The model receives a positive length reward whenever it generates a correct response that is shorter than the average of the previous correct ones. We denote this as HAPO w/ $h_{i}=mean$. Second, to ablate historical information from our length reward, we set $h_{i}$ to the average length of correct responses in the current batch. This effectively removes historical information and renders our method similar to Query-Opt (arora2025), except that we use our novel length reward. We denote this as HAPO w/o $h_{i}$. The results are shown in Table 5.
@@
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO w/ $h_{i}=min$ 79.08 661 81.05 2978 26.56 8171 62.23 3937
HAPO w/ $h_{i}=mean$ 74.46 413 74.90 2676 18.33 7487 55.90 3523
HAPO w/o $h_{i}$ 75.85 513 77.95 2754 24.60 11077 59.47 4781
Table 5: Results comparing different types of $h_{i}$.
HAPO w/ $h_{i}=mean$
Under this new update schedule, we expect $h_{i}$ to decrease more slowly during training, implying a slower reduction in response length. However, the results suggest otherwise. With the same number of training epochs, HAPO with the mean update produces more concise responses but lower accuracy than HAPO with the minimum update across the benchmarks. This indicates that HAPO tends to reduce response length more aggressively when using the historical mean. A possible explanation lies in the changes in reward distribution. When the minimum function is used, the lengths of most newly sampled responses are likely to exceed $h_{i}$, making many length rewards concentrate in the negative flat region. This results in more cases with a reward variance of 0, which consequently gives weaker learning signals. In contrast, using the mean function allows the sampled lengths to be more symmetrically distributed around $h_{i}$, thus leading to stronger learning signals.
HAPO w/o $h_{i}$
Compared with the original HAPO method, removing historical information increases the response length by a significant 36% on AIME2024, while slightly reducing the length on GSM8K and MATH500. The accuracy decreases across the three benchmarks. These results show that historical information is essential for our method to balance accuracy and efficiency.
Repetition Reward
GSM8K MATH500 AIME2024 Average
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$
DeepSeek-R1-1.5B 83.65 2066 79.15 6449 29.38 14615 64.06 7710
HAPO 79.08 661 81.05 2978 26.56 8171 62.23 3937
w/ Repetition Penalty 73.57 404 73.05 2782 17.71 6996 54.78 3394
Table 6: Results with repetition reward.
Following yeo2025, we incorporate an additional repetition penalty into the reward function. Specifically, we identify repeated n-grams within each generated response, count the total number of tokens that appear in these repeated n-grams, and divide this count by the total number of tokens in the response. The resulting repetition penalty, denoted as $rp$, is scaled by a weight constant $w_{\text{rp}}$ and subtracted from the overall reward. The final reward is given by:
Compared to HAPO, incorporating the repetition penalty results in shorter responses but leads to additional accuracy drops of approximately 5%, 8%, and 9% on GSM8K, MATH500, and AIME2024, respectively. We suspect that the penalty may be too strong. Using larger n-grams and reducing the weight $w_{\text{rp}}$ could potentially improve the overall performance.
Results with PPO
In addition to GRPO, we also experiment with PPO (schulam2017ppo) as the training algorithm. We set the learning rate to 3e-6 and train the models for 5 epochs. The results obtained using PPO are presented in Table 7.
GSM8K MATH500 AIME2024 Average
Models Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$ Pass@1 $\uparrow$ #Tokens $\downarrow$
@@
Discussion
Limitations and Future Work